Solve the system of equations:

y= x2 - 6x +3


y= -2x -6


1
Expert's answer
2021-02-25T05:04:42-0500

y=x2-6x+3

y=-2x-6

We equate the two equations as below;

x2-6x+3=-2x-6

Rearranging the above, we have;

x2-4x+9=0

We now solve the quadratic equation above.

The quadratic formula to be used is as below;

x=-b+- sqrt(b2-4ac)/2a

so;

x=-(-4)+- sqrt(-42-4*1*9)/2*1

x=4)+- sqrt(-20)/2

x = 2 - 2.24i

or

x = 2 + 2.24i


Now using the above values of x to find the possible values of y using the second equation.

y=-2x-6

y=-2(2 - 2.24i)-6

y=-10 +4.48i

or

y=-2x-6

y=-2(2 + 2.24i)-6

y=-10 -4.48i